OpenStudy (anonymous):

For each quadratic function, find the axis of symmetry of its graph. y = 3x^2 - 2x - 6

OpenStudy (unklerhaukus):

what is the graph doing at x=0

OpenStudy (anonymous):

thats the only question srry so idk

OpenStudy (unklerhaukus):

try and see \[y = 3x^2 - 2x - 6 \] when x=0 \[y=3\times 0^2-2\times0-6\]

OpenStudy (anonymous):

so then would it be -18. y=-18

OpenStudy (unklerhaukus):

0^2=0

OpenStudy (unklerhaukus):

so when x=0, y=-6

OpenStudy (anonymous):

so can i write it like this (0,-6)

OpenStudy (unklerhaukus):

yeh but that only a point in the parabola,

OpenStudy (unklerhaukus):

what is y when x>0, x<0

OpenStudy (anonymous):

for some reason i put (0,-6) and I got it wrong

OpenStudy (unklerhaukus):

well the axis of symmetry is a line and (0,-6) is just a point

OpenStudy (anonymous):

so then thats not the answer

OpenStudy (unklerhaukus):

If you look at the equation \[y = 3x^2 - 2x - 6 \] and then take the derivative with respect to x\[\frac{\text dy}{\text d x}= 36x - 2 \] you will get dy/dx=0 when x=1/18, dy/dx =0 mean you at at a turning point on the line, since a parabola only has one turning point this is your x value of the axis of symetry